OpenStudy (anonymous):

length of the segment with the folloowing enpoints G(-3,2) +(3,-3)

OpenStudy (anonymous):

length =\[\sqrt{(-3-3)^{2}+(-3-2)^{2}}\]

OpenStudy (mathmale):

You may find this problem easier to visualize and thus easier to solve if you plot these two points and then draw a line connecting them. It's the length of that line that you want; that's the distance between the two points. for starters: you could use either the "distance formula" or the "Pythagorean Theorem" (which are very much related to one another, by the way). Have you used the Draw utility (below) before?

OpenStudy (mathmale):

saketh30 is applying the Pythagorean Theorem here.

OpenStudy (anonymous):

length=\[\sqrt{36+25)}=\sqrt{61}\]

OpenStudy (mathmale):



OpenStudy (anonymous):

no i m applying the general formula to find the distance between two points (X1,Y1) and (x2,Y2)=\[\sqrt{(X2-X1)^{2}+(Y2-Y1)^{2}}\]